# Mesa-Extrapolation: A Weave Position Encoding Method for Enhanced Extrapolation in LLMs

### v1

@@ B.2 Params Setting
Implement Stair PE on RoPE and ALiBi   The handling of RoPE Su et al. (2023) involves assigning positions to the query vector and key vector. The length of the query corresponds to the length of the last chunk, while the lengths of the key and value align with the combined lengths of all chunks. We assign position encoding to the query vector according the same sequence positions. Regarding the key’s position encoding, we utilize the extrapolated width $E$ and extrapolated point $N$. By designing the position encoding of query and key respectively, the relative position scheme defined by Stair PE can be implemented. And, for computational efficiency, we ensure that the relative PE of the last token in the last chunk completely follows Stair PE. The other tokens of the last chunk are similar to Stair PE.
For ALiBi’s Press et al. (2021) position processing, we directly transform the relative position of the last chunk into the aforementioned structure within the mask matrix, refer to RoPE.
It is emphasized that the considerations outlined above primarily address the generation mechanism of LLM. The position encoding of the last chunk must align as closely as possible with the original position setting. This consideration is rooted in the belief that the latent concept Xie et al. (2021) formation heavily relies on there, with subsequent chunks providing essential information support. Experiments affirm that the last chunk significantly influences the accurate output of LLMs. This serves as the experimental foundation for our Mesa-Extrapolation method design.
ReRoPE & Leaky-ReRoPE    We follow Su (2023b) for configuring the parameters of ReRoPE and Leaky-ReRoPE. The extrapolated position for ReRoPE is established at 512, mirroring the setting for Leaky-ReRoPE. Simultaneously, the extrapolated increment for Leaky-ReRoPE is calculated in relation to the input length, following the formula below:
@@
LM-Infinite & Streaming-LLM   LM-Infinite Han et al. (2023) introduces two branches for attention masking: a global branch on the left and a local branch on the right. The global branch enables each token to attend to the preceding $n_{global}$ tokens if they appear before the current token. Conversely, the local branch allows each token to attend to preceding tokens within a distance of $n_{local}$. Tokens outside these two branches are disregarded during the attention operation. Subsequently, following its setting, set $n_{local}=\mathcal{I}$, where $\mathcal{I}$ still represents the training length. The choice of $n_{global}$ has a minor impact on model performance within the range $[10,100]$, and thus, we fix $n_{global}=100$ as its code setting. The distance limit entails the "effective distance" $d$ within $\mathcal{I}$, and we set $d=\mathcal{I}$.
Streaming-LLM Xiao et al. (2023) adopts a similar design to LM-Infinite. It introduces an initial token known as Attention Sinks, whose length is denoted by $x$. It also defines a Rolling KV Cache for retaining the most recent tokens, whose length is denoted by $y$. It set $x=4$ and $y=\mathcal{T}-x$, where $\mathcal{T}$ denotes the maximum training length. The central tokens are referred to Evicted Tokens.
@@
Origin & Dynamic-NTK   For Origin, it refers to loading the model directly, without using any methods. Dynamic-NTK differs from Origin by only adjusting the angle value of its RoPE component, according to the input length Liu et al. (2023).
@@ B.3 Weave PE-based Schemes
While these weave PE schemes, including ReRoPE, Leaky-ReRoPE, and Stair PE, demonstrate theoretical feasibility, their practical implementation may encounter computational complexities. For instance, both ReRoPE and Leaky-ReRoPE are primarily employed in the LLaMA models, which is based on RoPE Su et al. (2023). Due to the distinctive nature of RoPE, these methods necessitate the calculation of the attention matrix more than once, resulting in double the normal memory consumption. Moreover, given the quadratic complexity of the calculations, as the input length expands, both inference time and memory consumption grow proportionally.
@@ C.1 BLEU Results on GovReport
We use the GovReport dataset and employ BLEU Papineni et al. (2002) as evaluation metric, with a same setting applied on Table 1. The experimental result for BLEU is presented in Table 2. In Table 2, we present the mean and standard variance for each method across various interval lengths. It is evident that, as the input token length increases, the performance consistently deteriorates for LM-Infinite. Streaming-LLM also experiences extrapolation failure beyond the model’s inherent effective window length of 4k.
@@ C.2 Perplexity (PPL) metrics on MPT-7b
ALiBi Press et al. (2021) can still effectively extrapolate to around $3.5$k position after surpassing the maximum training length of $2$k. This observation affirms the extrapolation capability of the ALiBi method, especially when compared with RoPE’s extrapolation results.
However, it is essential to highlight that while ALiBi achieves extrapolation to $3.5$k in above NLL task, its performance in the Passkey and LongEval tasks (refer to C.3) reveals limitations. This discrepancy is attributed to an inherent characteristic of ALiBi. ALiBi’s expression (refer to D.2) indicates a constant decrease in the value of its position encoding as the relative distance increases. Considering the $\mathrm{SoftMax}$ operation within the attention mechanism, under identical conditions, the attention score obtained becomes smaller. This implies that tokens situated farther away receive less attention, possibly leading to the neglect of these distant tokens. Therefore, despite ALiBi’s apparent success in effectively extrapolating the NLL task to $3.5$k, it falls short in attending to tokens beyond the maximum training length in practice. A similar analysis process is as Su (2023a).
It’s noted that, NoPE Kazemnejad et al. (2023) claims to possess extrapolation capabilities akin to ALiBi. We hypothesize that the underlying reasons for NoPE’s extrapolation performance align with the analysis presented here.
@@ C.3 Evaluation on LongEval
We conduct additional testing on LongEval Krishna et al. (2023) lines task, a recently prominent evaluation task for long texts.
@@ C.4 Evaluation on LongBench
We select LongBench Bai et al. (2023) dataset and use $5$ major categories of tasks, including Single-Document QA, Multi-Document QA, Few-shot Learning, Synthesis Tasks and Code Completion. Among them, each task selects a dataset, namely qasper, hotpotqa, samsum, passage-retrieval-en, and repobench-p. Taking into account the varying memory consumption of different methods, we opt to filter out samples with an input length exceeding $10$k to prevent discrepancies caused by out-of-memory (OOM) issues. We use the abbreviations in Table 3: S-Document (Single Document) QA, M-Document (Multi Document) QA, F-Learning (Few-shot Learning), S-Tasks (Synthesis Tasks), C-Completion (Code Completion).
@@ C.5 Enhanced Prompts for Mesa-Extrapolation
Based on these considerations, we employ the model’s corresponding prompt Geng, Xinyang and Liu, Hao (2023) and augment it with explicit goal statements, referred to as an enhanced prompt, like below:
@@
Additionally, we leverage model parallel inference technology BlackSamorez (2023) and perform verification concurrently on $2$ A800 GPUs.
@@ C.7 Theoritical Speed & Memory
Table 4 shows the theoretical result about the memory usage. For Origin, the attention matrix calculation requires the current token to compute the attention score with each previous token, resulting in a memory footprint of $\mathcal{O}(n^{2})$. Dynamic-NTK only alters the angle base, making it different from the Origin but still quadratic. For ReRoPE and Leaky-ReRoPE, as showed in Su (2023b), since the attention matrix needs to be calculated twice, their memory footprint is $2\times\mathcal{O}(n^{2})$. For Mesa-Extrapolation, its strategy involves splitting chunks while avoiding the quadratic term. Accounting for splicing in the last chunk, its total memory usage scales proportional to $\mathcal{O}((2+\sqrt{2})n)$. LM-Infinite and Streaming-LLM adopt a similar $\Lambda$-shaped mask, that scales proportional to $\mathcal{O}((1+\sqrt{2})n)$.
@@ C.8 Generated Summary using Mesa-Extrapolation
We present here the results of the summary generation task using GovReport Huang et al. (2021). In this experiment, task is to generate a summary for texts of varying lengths, limited to 1000 tokens. The outcomes yielded by Mesa-Extrapolation are outlined below on Table LABEL:tab:flency-mesa:
@@ C.9 Evaluation of Phi-3-mini-128k-instruct Model on Ruler Datasets
We further conducted experimental validation on the Ruler datasets Hsieh et al. (2024), focusing on the single-keys NIAH task. The needle-in-a-haystack (NIAH) test assesses the ability to retrieve a specific piece of information (the “needle”) from long distractor texts (the “haystack”).
@@ D.1 Preliminary
In this section, we lay the groundwork and introduce the notation for below theorem analysis. Bold letters typically denote vectors or matrices, while normal letters represent scalars. We mainly follow Kazemnejad et al. (2023) to define these notations.
@@
where $\sigma$ is a non-linear activation function(usually $\mathrm{ReLU}$ or $\mathrm{GeLU}$ Hendrycks & Gimpel (2016)), and $\sigma(\cdot)\in\mathbb{R}^{d}$ is layer normalization Ba et al. (2016). It’s worthy noted that the additive view of attention heads Elhage et al. (2021) in Eq.(5) is mathematically equivalent with the view of concatenate and multiple Vaswani et al. (2017). The additive view is easier to understand and analyze.
@@ D.2 NoPE & PE
For ALiBi Press et al. (2021), it takes the form:
@@
For RoPE Su et al. (2023), it’s a relative PE that applies a rotation to the query and key representations based on their absolute positions before dot product attention. We formulate RoPE for model dimension $d=2$, and its dot product as below:
@@ Appendix E Proofs
We provide proof about Theorems 3.1, 3.2, 3.3, and Corollary 4.1. Our proofs are inspired by Kazemnejad et al. (2023) Weiss et al. (2021) Lindner et al. (2023) and relies on the causal attention mask in the decoder-only Transformer and the SoftMax function.
@@
We conjecture that while the attention mechanism in RoPE Su et al. (2023) involves multiplying by the rotation matrix of the relative position, it’s essential to acknowledge that neural networks possess the capability to model any function. This implies the potential to decompose the relative position information into an additive form, as exemplified in our extrapolation scheme detailed in Theorem 3.3.
@@ Appendix G Stair PE and Self-Extend
According to its equation $3$ in Self-Extend (Jin et al. (2024)), $P=P\mathbin{//}G$, and the shifted relative position $(W-W\mathbin{//}G)$, we can get the position of query as:

### v3

@@ B.2 Params Setting
Implement Stair PE on RoPE and ALiBi   The handling of RoPE [36] involves assigning positions to the query vector and key vector. The length of the query corresponds to the length of the last chunk, while the lengths of the key and value align with the combined lengths of all chunks. We assign position encoding to the query vector according the same sequence positions. Regarding the key’s position encoding, we utilize the extrapolated width $E$ and extrapolated point $N$. By designing the position encoding of query and key respectively, the relative position scheme defined by Stair PE can be implemented. And, for computational efficiency, we ensure that the relative PE of the last token in the last chunk completely follows Stair PE. The other tokens of the last chunk are similar to Stair PE.
For ALiBi’s [29] position processing, we directly transform the relative position of the last chunk into the aforementioned structure within the mask matrix, refer to RoPE.
It is emphasized that the considerations outlined above primarily address the generation mechanism of LLM. The position encoding of the last chunk must align as closely as possible with the original position setting. This consideration is rooted in the belief that the latent concept [46] formation heavily relies on there, with subsequent chunks providing essential information support. Experiments affirm that the last chunk significantly influences the accurate output of LLMs. This serves as the experimental foundation for our Mesa-Extrapolation method design.
ReRoPE & Leaky-ReRoPE    We follow [35] for configuring the parameters of ReRoPE and Leaky-ReRoPE. The extrapolated position for ReRoPE is established at 512, mirroring the setting for Leaky-ReRoPE. Simultaneously, the extrapolated increment for Leaky-ReRoPE is calculated in relation to the input length, following the formula below:
@@
LM-Infinite & Streaming-LLM   LM-Infinite [14] introduces two branches for attention masking: a global branch on the left and a local branch on the right. The global branch enables each token to attend to the preceding $n_{global}$ tokens if they appear before the current token. Conversely, the local branch allows each token to attend to preceding tokens within a distance of $n_{local}$. Tokens outside these two branches are disregarded during the attention operation. Subsequently, following its setting, set $n_{local}=\mathcal{I}$, where $\mathcal{I}$ still represents the training length. The choice of $n_{global}$ has a minor impact on model performance within the range $[10,100]$, and thus, we fix $n_{global}=100$ as its code setting. The distance limit entails the "effective distance" $d$ within $\mathcal{I}$, and we set $d=\mathcal{I}$.
Streaming-LLM [45] adopts a similar design to LM-Infinite. It introduces an initial token known as Attention Sinks, whose length is denoted by $x$. It also defines a Rolling KV Cache for retaining the most recent tokens, whose length is denoted by $y$. It set $x=4$ and $y=\mathcal{T}-x$, where $\mathcal{T}$ denotes the maximum training length. The central tokens are referred to Evicted Tokens.
@@
Origin & Dynamic-NTK   For Origin, it refers to loading the model directly, without using any methods. Dynamic-NTK differs from Origin by only adjusting the angle value of its RoPE component, according to the input length [24].
@@ B.3 Weave PE-based Schemes
While these weave PE schemes, including ReRoPE, Leaky-ReRoPE, and Stair PE, demonstrate theoretical feasibility, their practical implementation may encounter computational complexities. For instance, both ReRoPE and Leaky-ReRoPE are primarily employed in the LLaMA models, which is based on RoPE [36]. Due to the distinctive nature of RoPE, these methods necessitate the calculation of the attention matrix more than once, resulting in double the normal memory consumption. Moreover, given the quadratic complexity of the calculations, as the input length expands, both inference time and memory consumption grow proportionally.
@@ C.1 BLEU Results on GovReport
We use the GovReport dataset and employ BLEU [27] as evaluation metric, with a same setting applied on Table 1. The experimental result for BLEU is presented in Table 2. In Table 2, we present the mean and standard variance for each method across various interval lengths. It is evident that, as the input token length increases, the performance consistently deteriorates for LM-Infinite. Streaming-LLM also experiences extrapolation failure beyond the model’s inherent effective window length of 4k.
@@ C.2 Perplexity (PPL) metrics on MPT-7b
ALiBi [29] can still effectively extrapolate to around $3.5$k position after surpassing the maximum training length of $2$k. This observation affirms the extrapolation capability of the ALiBi method, especially when compared with RoPE’s extrapolation results.
However, it is essential to highlight that while ALiBi achieves extrapolation to $3.5$k in above NLL task, its performance in the Passkey and LongEval tasks (refer to C.3) reveals limitations. This discrepancy is attributed to an inherent characteristic of ALiBi. ALiBi’s expression (refer to D.2) indicates a constant decrease in the value of its position encoding as the relative distance increases. Considering the $\mathrm{SoftMax}$ operation within the attention mechanism, under identical conditions, the attention score obtained becomes smaller. This implies that tokens situated farther away receive less attention, possibly leading to the neglect of these distant tokens. Therefore, despite ALiBi’s apparent success in effectively extrapolating the NLL task to $3.5$k, it falls short in attending to tokens beyond the maximum training length in practice. A similar analysis process is as [34].
It’s noted that, NoPE [20] claims to possess extrapolation capabilities akin to ALiBi. We hypothesize that the underlying reasons for NoPE’s extrapolation performance align with the analysis presented here.
@@ C.3 Evaluation on LongEval
We conduct additional testing on LongEval [21] lines task, a recently prominent evaluation task for long texts.
@@ C.4 Evaluation on LongBench
We select LongBench [3] dataset and use $5$ major categories of tasks, including Single-Document QA, Multi-Document QA, Few-shot Learning, Synthesis Tasks and Code Completion. Among them, each task selects a dataset, namely qasper, hotpotqa, samsum, passage-retrieval-en, and repobench-p. Taking into account the varying memory consumption of different methods, we opt to filter out samples with an input length exceeding $10$k to prevent discrepancies caused by out-of-memory (OOM) issues. We use the abbreviations in Table 3: S-Document (Single Document) QA, M-Document (Multi Document) QA, F-Learning (Few-shot Learning), S-Tasks (Synthesis Tasks), C-Completion (Code Completion).
@@ C.5 Enhanced Prompts for Mesa-Extrapolation
Based on these considerations, we employ the model’s corresponding prompt [13] and augment it with explicit goal statements, referred to as an enhanced prompt, like below:
@@
Additionally, we leverage model parallel inference technology [6] and perform verification concurrently on $2$ A800 GPUs.
@@ C.7 Theoritical Speed & Memory
Table 4 shows the theoretical result about the memory usage. For Origin, the attention matrix calculation requires the current token to compute the attention score with each previous token, resulting in a memory footprint of $\mathcal{O}(n^{2})$. Dynamic-NTK only alters the angle base, making it different from the Origin but still quadratic. For ReRoPE and Leaky-ReRoPE, as showed in [35], since the attention matrix needs to be calculated twice, their memory footprint is $2\times\mathcal{O}(n^{2})$. For Mesa-Extrapolation, its strategy involves splitting chunks while avoiding the quadratic term. Accounting for splicing in the last chunk, its total memory usage scales proportional to $\mathcal{O}((2+\sqrt{2})n)$. LM-Infinite and Streaming-LLM adopt a similar $\Lambda$-shaped mask, that scales proportional to $\mathcal{O}((1+\sqrt{2})n)$.
@@ C.8 Generated Summary using Mesa-Extrapolation
We present here the results of the summary generation task using GovReport [18]. In this experiment, task is to generate a summary for texts of varying lengths, limited to 1000 tokens. The outcomes yielded by Mesa-Extrapolation are outlined below on Table LABEL:tab:flency-mesa:
@@ C.9 Evaluation of Phi-3-mini-128k-instruct Model on Ruler Datasets
We further conducted experimental validation on the Ruler datasets [17], focusing on the single-keys NIAH task. The needle-in-a-haystack (NIAH) test assesses the ability to retrieve a specific piece of information (the “needle”) from long distractor texts (the “haystack”).
@@ D.1 Preliminary
In this section, we lay the groundwork and introduce the notation for below theorem analysis. Bold letters typically denote vectors or matrices, while normal letters represent scalars. We mainly follow [20] to define these notations.
@@
where $\sigma$ is a non-linear activation function(usually $\mathrm{ReLU}$ or $\mathrm{GeLU}$ [16]), and $\sigma(\cdot)\in\mathbb{R}^{d}$ is layer normalization [2]. It’s worthy noted that the additive view of attention heads [11] in Eq.(5) is mathematically equivalent with the view of concatenate and multiple [41]. The additive view is easier to understand and analyze.
@@ D.2 NoPE & PE
For ALiBi [29], it takes the form:
@@
For RoPE [36], it’s a relative PE that applies a rotation to the query and key representations based on their absolute positions before dot product attention. We formulate RoPE for model dimension $d=2$, and its dot product as below:
@@ Appendix E Proofs
We provide proof about Theorems 3.1, 3.2, 3.3, and Corollary 4.1. Our proofs are inspired by [20] [43] [23] and relies on the causal attention mask in the decoder-only Transformer and the SoftMax function.
@@
We conjecture that while the attention mechanism in RoPE [36] involves multiplying by the rotation matrix of the relative position, it’s essential to acknowledge that neural networks possess the capability to model any function. This implies the potential to decompose the relative position information into an additive form, as exemplified in our extrapolation scheme detailed in Theorem 3.3.
@@ Appendix G Stair PE and Self-Extend
According to its equation $3$ in Self-Extend ([19]), $P=P\mathbin{//}G$, and the shifted relative position $(W-W\mathbin{//}G)$, we can get the position of query as: